OpenStudy (anonymous):

Solve for f. f + (–15) = –32 A. 17 B. 47 C. –17 D. –47

OpenStudy (anonymous):

C

OpenStudy (anonymous):

f - 15 = -32 f = -32 +15 f = -17

OpenStudy (anonymous):

So Your answer Is c

OpenStudy (shamim):